# thanks: Contributed equallythanks: Contributed equally

### v2

@@ VI Conclusions
To conclude, we present a novel mechanism for formation of fluid flocks that only requires long-ranged (net) repulsive torques and short-ranged repulsive forces from steric interactions. We have shown that this is equivalently abstracted to a sliding-whilst-rotating mechanism at the pair colloidal level. Although the roles of excluded volume interactions [19, 21, 15, 16], repulsive torques - either short [15] or long [26] ranged - and long-ranged repulsive forces [21, 26] have been studied separately, we have analyzed here the individual contributions of each of those components, and argued that the former two are the minimal requisite components. Our mechanism can summarized as follows: the pair colloids slide together for at least a unit length before deterministically rotating away from each other. This is required to break the forward-backward symmetry at the pair-collision level. Destabilizing of the flock arises due to symmetric collisions at the pair colloidal level, wherein they slide exceedingly short before rotating away upon collision. These results we show to be consistent with a coarse-grained hydrodynamic theory, though the transition lines are greatly overestimated wherein, indicating the need to go beyond conventional coarse-graining hypotheses [17, 16]. We then use this sliding-whilst-rotating mechanism as a basis for a phenomenological agent-based model based on update rules of colliding pair colloids, where we further find that it reproduces our results well with only one additional free parameter; further substantiating the underlying fluid flock forming mechanism.
@@
The CLF (or the polar liquids) are both generated by, and further destabilized by, the deterministic torques themselves. Notably, the apparent ubiquitous role of noise in destabilizing flocks [4, 11, 21, 26] is also absent in our mechanism (these are negligibly small in our simulation). We note that in the absence of short-ranged excluded volume repulsions, our results would converge to that recently studied in [26] where the need for (long-ranged) repulsive forces is more imminent. Translational repulsion merely renders the flock to acquire a crystalline structure. We also have not studied here the role of particle shape anisotropy [17, 14], which would add additional competing length scales to our analysis. Finally, the agent-based model presented here being minimal, and still does not account for the crystalline flock for $l_{t}>0$ (due to purely local interaction rules), which suggests extensions for future work. We expect our results here to be direcstly relevant to a variety of experimental systems. For example, migrating cell collections have been widely reported to exhibit spatial organization resembling a polar fluid [28, 22], while their interaction mechanism has been known to include (among others) long-ranged chemotaxis and avoidance torques [27]. The presented mechanism being very generic, we expect it to also be relevant for interacting colloids of non-chemical origin [25, 38, 37], where the underlying mechanism should also hold.
@@ Appendix A Simulation details
FIG $L$ $dt$ $b$ $v_{s}$ $\kappa$ $N$ $\chi_{t}$ $\chi_{r}$ $D_{r}$ $D_{t}$
1 $100$ $0.01$ $1$ $50$ $175$ $(319,2578)$ 00 $(0,7)$ $10^{-6}$ $10^{-8}$
3.(a) $100$ $0.01$ $1$ $50$ $175$ $1024$ $(0,5)$ $(0,10)$ $10^{-6}$ $10^{-8}$
3.(b) $100$ $0.01$ $1$ $50$ $175$ $1024$ $(-1,0)$ $(0,10)$ $10^{-6}$ $10^{-8}$
3.(d) $186$ $0.01$ $1$ $50$ $175$ $4000$ 00 $0.75$ $10^{-6}$ $10^{-8}$
3.(e) $186$ $0.01$ $1$ $50$ $175$ $4000$ $3$ $6$ $10^{-6}$ $10^{-8}$
4.(a) $(77,354)$ $0.01$ $1$ $50$ $175$ $1600$ 00 $(0,9)$ $10^{-6}$ $10^{-8}$
4.(b) $118$ $0.01$ $1$ $50$ $175$ $1600$ $(0,12)$ $(0,32)$ $10^{-6}$ $10^{-8}$
5.(b) $100$ $0.1$ NA $5$ NA $(319,2578)$ 00 $(0,15)$ NA NA
Table 1: Parameter values used for respective figures in the main text. Note that we consider noise to be less dominant to deterministic effects. Indeed, for the results in this paper, the Péclet number $\mathrm{Pe}=v_{s}/(bD_{r})$ is around $10^{9}$.
The update equations is given in Eq. 1 of the main text. The position and orientations of the particles are updated using a forward Euler-Maruyama method. Initially particles are distributed randomly over the two-dimensional space. The initial orientations are also randomly distributed over the range of angles [$-\pi,+\pi$] (angles are computed with respect to the positive $x$-axis). Periodic boundary conditions are applied on both the $x$-axis and $y$-axis. To compute chemical interactions using the expression of $\mathbf{J}_{i}$, we need to sum over periodic images. We have used a regular summation convention (using minimum image convention) for all results reported here, and checked these with the full Ewald summation convention for selected parameter values. The parameter values used are given here: radius of a colloid $b=1$, velocity $v_{s}=50$, total number of particles $N$ is variable according to the figure. All simulations are performed using periodic boundary conditions with a regular summation convention. A summary of all of these is given in the Table (1). For the agent-based model (Fig.5), additional parameters used are mentioned in the caption of Fig.5.
@@ Appendix B Correlation functions and fluctuations
An example of this for isotropic phase is given in Fig. 6(a). The pair distribution functions are determined for the four different $\theta$ ranges as: $\theta$:[$20,\pi/$ 2], $\theta$:[$\pi/2,\pi$], $\theta$:[$-\pi,-\pi/2$] and $\theta$:[$-\pi/2,0$]. We see that in the isotropic phase, the probability of finding a neighboring particle is different in the front relative to the back of the direction of motion (orientation), indicative of the symmetry breaking in an active disordered gas [17, 38, 26]. In the main text, we show that $g(r,\phi)$ also distinguishes CLFs from CCFs, the former having a more quadrant-like structure compared to the latter’s hexatic structure. In addition, we compare the structure of $g(r,\phi)$ for the band phases in Fig. (6)(b), for the CLF and CCF density bands. Note that the band structure is more dense in the case of $\chi_{t}=0$, resulting a more intense correlation in $g(r,\phi)$. A related quantity that we compute is the radial $g(r)$, given by
@@
this is plotted in Fig. (3)(f), distinguishing the CLF and CCF.
To further compare the CLF and CCF, we measure the following quantities.
B.1 Velocity correlations
We define the velocity correlation as
$C_{v}(r)=\langle\mathbf{e}_{i}\mathbf{e}_{j}\rangle_{|\mathbf{r}_{i}-\mathbf{r}_{j}|-r}.$
The above is plotted for the two cases of CLF and CCF in Fig.7.
B.2 Orientational (angular) fluctuations
[…]
B.3 Giant number fluctuations.
[…]
Figure 7: Velocity correlations for the CLF (left) and CCF (right). In both cases, the disordered gas shows exponential decay to zero at large distances. In the flocking case, both are unity across long distances (globally polarized). In the band phase (green), the correlations in the CLF decay whereas those in the CCF do not.
Figure 8: Comparison of CLF and CCF via (a) density fluctuations, (b) orientational fluctuations, and (c) giant number fluctuations. For the CLF, (a) shows residual fluctuations, whereas for the CCF, the fluctuations die out in the steady-state. In (b), a similar phenomenon is seen for the orientational fluctuations. Note that both curves are strongly correlated (shared monotonicity). In (c) the giant number fluctuations of both show an enhanced exponent $\sim 1.2$.
@@ Appendix C Description of the supplementary movies
  • Movie I. In the case of $\chi_{t}=0$ (CLF), the liquid flock gradually forms throughout the system and moves in a particular direction. The fixed parameter values are: $L=418$, $b=1$, $dt=0.01$, $D_{r}=10^{-3}$, $D_{t}=10^{-3}$, $\kappa=175$, $v_{s}=50$, $\chi_{t}=0$, $\chi_{r}=0.75$.
  • Movie II. In the case of $\chi_{t}>0$ (CCF), the particles first start to form local polar flocks (colored patch), then merge and pick a particular direction of motion entirely. The fixed parameter values are: $L=418$, $b=1$, $dt=0.01$, $D_{r}=10^{-3}$, $D_{t}=10^{-3}$, $\kappa=175$, $v_{s}=50$, $\chi_{t}=3$, $\chi_{r}=6$.

### v3

@@ V Conclusions
To conclude, we present a minimal mechanism for formation of fluid flocks that only requires long-ranged (net) repulsive torques and short-ranged repulsive forces from steric interactions. Although the roles of excluded volume interactions [19, 21, 15, 16], repulsive torques - either short [15] or long [26] ranged - and long-ranged repulsive forces [21, 26] have been studied separately, we have analyzed here the individual contributions of each of those components, and argued that the former two are the minimal requisite components. Our mechanism can summarized as follows: the pair colloids slide together for at least a unit length before deterministically rotating away from each other. This is required to break the forward-backward symmetry at the pair-collision level. Destabilizing of the flock arises due to symmetric collisions at the pair colloidal level, wherein they slide exceedingly short before rotating away upon collision. These results we show to be consistent with a minimal hydrodynamic theory. One may ask whether such a sliding mechanism alone added to cognitive based (agent based) models [10, 11] could reproduce the flocking transition; indeed we will leave this to future work.
@@
The CLF (or the polar liquids) are both generated by, and further destabilized by, the deterministic torques themselves. Notably, the apparent ubiquitous role of noise in destabilizing flocks [4, 10, 21, 26] is also absent in our mechanism (these are negligibly small in our simulation). We note that in the absence of short-ranged excluded volume repulsions, our results would converge to that recently studied in [26] where the need for (long-ranged) repulsive forces is more imminent. Translational repulsion merely renders the flock to acquire a crystalline structure. We also have not studied here the role of particle shape anisotropy [17, 14], which would add additional competing length scales to our analysis. We expect our results here to be directly relevant to a variety of experimental systems. For example, migrating cell collections have been widely reported to exhibit spatial organization resembling a polar fluid [28, 22], while their interaction mechanism has been known to include (among others) long-ranged chemotaxis and avoidance torques [27]. The presented mechanism being very generic, we expect it to also be relevant for interacting colloids of non-chemical origin [25, 38, 37], where the underlying mechanism should also hold.
Conflicts of interest
There are no conflicts to declare.
Data availability
The datasets are generated from simulation. They are available from the corresponding authors on reasonable request.
@@ Appendix A Simulation details
FIG $L$ dtdt $b$ $v_{0}$ $\kappa$ NN $\chi_{t}$ $\chi_{r}$ $D_{r}$ $D_{t}$
1 $100$ $0.01$ 11 $50$ $175$ $(319,2578)$ 0 $(0,7)$ $10^{-6}$ $10^{-8}$
3.(a) $100$ $0.01$ 11 $50$ $175$ $1024$ $(0,5)$ $(0,10)$ $10^{-6}$ $10^{-8}$
3.(b) $100$ $0.01$ 11 $50$ $175$ $1024$ (1,0)(-1,0) $(0,10)$ $10^{-6}$ $10^{-8}$
3.(d) $186$ $0.01$ 11 $50$ $175$ $4000$ 0 $0.75$ $10^{-6}$ $10^{-8}$
3.(e) $186$ $0.01$ 11 $50$ $175$ $4000$ 33 66 $10^{-6}$ $10^{-8}$
4.(a) $(77,354)$ $0.01$ 11 $50$ $175$ $1600$ 0 $(0,9)$ $10^{-6}$ $10^{-8}$
4.(b) $118$ $0.01$ 11 $50$ $175$ $1600$ $(0,12)$ $(0,32)$ $10^{-6}$ $10^{-8}$
5.(a) $100$ $0.01$ 11 $50$ $175$ $(637,1592)$ 0 $(0,8)$ $10^{-6}$ $10^{-8}$
5.(b) $100$ $0.01$ 11 $50$ $175$ $(637,1592)$ $(0,10)$ $27$ $10^{-6}$ $10^{-8}$
Table 1: Parameter values used for respective figures of the paper. Here $L$ is system size, $b$ is the radius of the particle, NN is the number of particles, dtdt is the time-stepping of the Euler-Maruyama integrator. The remaining parameters are defined after Eq.(1). Note that we consider noise to be less dominant to deterministic effects. Indeed, for the results in this paper, the Péclet number $\mathrm{Pe}=v_{0}/(bD_{r})$ is around $10^{9}$.
The update equations is given in Eq. 1 of the main text. The position and orientations of the particles are updated using a forward Euler-Maruyama method. Initially particles are distributed randomly over the two-dimensional space. The initial orientations are also randomly distributed over the range of angles $\pi,+\pi$] (angles are computed with respect to the positive $x$-axis). Periodic boundary conditions are applied on both the $x$-axis and $y$-axis. To compute chemical interactions using the expression of $\mathbf{J}_{i}$, we need to sum over periodic images. We have used a regular summation convention (using minimum image convention) for all results reported here, and checked these with the full Ewald summation convention for selected parameter values. A summary of all of the values of parameters used for the simulation are given in Table (1).
@@ Appendix B Correlation functions and fluctuations
An example of this for isotropic phase is given in Fig. 6(a). The pair distribution functions are determined for the four different $\theta$ ranges as: $\theta$$0,\pi/2$], $\theta$$\pi/2,\pi$], $\theta$$\pi,-\pi/2$] and $\theta$$\pi/2,0$]. We see that in the isotropic phase, the probability of finding a neighboring particle is different in the front relative to the back of the direction of motion (orientation), indicative of the symmetry breaking in an active disordered gas [17, 38, 26]. In the main text, we show that $g(r,\phi)$ also distinguishes CLFs from CCFs, the former having a more quadrant-like structure compared to the latter’s hexatic structure. In addition, we compare the structure of $g(r,\phi)$ for the band phases in Fig. (6)(b), for the CLF and CCF density bands. Note that the band structure is more dense in the case of $\chi_{t}=0$, resulting a more intense correlation in $g(r,\phi)$. A related quantity that we compute is the radial $g(r)$, given by
@@
this is plotted in Fig. (3), distinguishing the CLF and CCF.
An additional feature of the CLF is that they have a steady-state persistent fluctuations of both density and polarization fields (dynamical steady-state - note that our simulations are effectively noiseless), which is absent in the CCF. In Fig.7, we show fluctuations for positions and orientations for CLF and CCF . Here, the Density fluctuations is given by: $\langle(\Delta\rho)^{2}\rangle=\langle\rho^{2}\rangle-\langle\rho\rangle^{2}$. Here, $\langle\quad\rangle$ denotes a spatial average. Note that the $\rho$ field has to be obtained via binning. We also measure Orientational (angular) fluctuations, which is given by $\langle(\Delta\theta)^{2}\rangle=\langle\theta^{2}\rangle-\langle\theta\rangle^{2}$. Here, $\langle\quad\rangle$ denotes average over colloids.
Figure 7: Comparison of CLF ($\chi_{t}=0$) and CCF ($\chi_{t}>0$) via density fluctuations ($\langle(\Delta\rho)^{2}\rangle$), and orientational fluctuations ($\langle(\Delta\theta)^{2}\rangle$). For the CLF, the panel on the left shows residual fluctuations, whereas for the CCF, the fluctuations die out in the steady-state. In the panel on the right, a similar phenomenon is seen for the orientational fluctuations. Note that both curves are strongly correlated (shared monotonicity).
@@ Appendix C Description of the supplementary movies
  • Movie I. In the case of $\chi_{t}=0$ (CLF), the liquid flock gradually forms throughout the system and moves in a particular direction. The fixed parameter values are: $L=418$, $b=1$, $dt=0.01$, $D_{r}=10^{-3}$, $D_{t}=10^{-3}$, $\kappa=175$, $v_{0}=50$, $\chi_{t}=0$, $\chi_{r}=0.75$.
  • Movie II. In the case of $\chi_{t}>0$ (CCF), the particles first start to form local polar flocks (colored patch), then merge and pick a particular direction of motion entirely. The fixed parameter values are: $L=418$, $b=1$, $dt=0.01$, $D_{r}=10^{-3}$, $D_{t}=10^{-3}$, $\kappa=175$, $v_{0}=50$, $\chi_{t}=3$, $\chi_{r}=6$.